### v3

# Further Comments on Yablo’s Construction ††thanks: File: WAB, [Sch25], arXiv 2504.10370

Abstract
We continue our analysis of Yablo’s coding of the liar paradox by infinite acyclic graphs, [Yab82].
The present notes are based on and continue the author’s previous results on the problem.
In particular, our approach is often more systematic than before.
Contents
1 Overview
These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions. All these papers are, of course, based on Yablo’s seminal [Yab82].
See also the author’s earlier versions on arXiv.
In Section 2 (page 2), we first present the background, its logical and graph theoretical side, and their interplay. We also discuss a useful third truth value, which expresses that neither nor is consistent.
In most cases, we will work with simple conjunctions (with and without negations).
In Section 3 (page 3), we discuss elementary contradiction (cells). Essentially, these are variations of Yablo’s triangles. The simplest contradictions are too simple, they have an “escape possibility”, i.e., combining them will leave one possibility without contradictions. Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework.
In Section 4 (page 4), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction, but give a finer analysis, discuss its properties, and again escape possibilities (in indirect composition).
We may start the construction with some simple but we will automatically end in a Yablo like construction, when satisfying the basic requirements (C1) and (C2) (see below, Section 1.1 (page 1.1)). This is discussed in the sections on “Saw Blades”, see Section 7.6 in [Sch22] and Section 3 in [Sch23b].
- (4)
1.1 Basic Terminology
Yablo’s construction
Yablo’s construction is an infinite directed graph with nodes and (negative) arrows for
Nodes stand for propositional variables, and the meaning of arrows is that variable occurs (negatively) in the formula attached to in most cases it expresses (We will identify for simplicity nodes with their attached variables.)
It is easy to see that no node may be true, nor false: Suppose is true, than all are false, so must be false, then contradiction, as all must be false. Suppose is false, then some must be true, again a contradiction (to the above).
Yablo Cell, Yablo Triangle
The basic contradiction structure has the form with the meaning (so and
We sometimes call the head, the knee, and the foot of the Yablo Cell (or Triangle).
Conditions (C1) and (C2)
We sometimes write when we want to emphasize that we assume is positive, similarly for the negative case.
Consider in the Yablo construction (but this applies to all
We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction. (More precisely, see Section 4.1 (page 4.1), (3).)
1.2 Comments On the Strategy
These comments will be clearer in hindsight, after having read the details. Still the best place to put them seems to be in the introduction.
The comments on “ is false ” show that there is no other way to express this in an acyclic graph of our type.
The comments on the construction in the limit solve a seeming contradiction by chosing the right level of abstraction, i.e. by arguing on the level of instances. We do not show that there cannot be any other way to do this - we do not know, we just see no other possibility.
1.2.1 Expressing “x is false” In an Acyclic Graph
is sometimes expressed by and by a contradiction.
How can we express that is wrong in a finite acyclic graph?
Consider the set of minimal elements below , and the formula corresponding to the graph in DNF of these elements. This might be
(the graph is trivial, just
or a tautology or contradiction.
a disjunction of conjunctions of elements in , or their negations. By acyclicity,
So the only way to express that a is wrong is by a contradiction.
1.2.2 Satisfying the Antagonistic Requirements (C1) and (C2) In the Limit
We want to express by an infinite graph that both and are inconsistent.
By (102.2), this must be expressed by suitable contradictions. Therefore, we treat such contradictions as basic building blocks, as Yablo did.
The problems are then
- (1)
How do we combine them to obtain a suitable graph expressing that both and are impossible?
The problem is that, if is contradictory, then is a tautology in classical logic, and thus in the theory expressed by a finite acyclic graph. So we alternate between and a tautology, and it is not clear what such constructions are in the limit, when we leave classical logic.
This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 4 (page 4).
We have to look at instances.
- (2.1)
- (2.2)
But is now a branching point. The branch continues to be satisfied for (C1), but the (new) branch opens a new problem for (C1), as is an OR for
- (2.4)
So we have new instances of problems for (C1), (C2), we solve the new instances, say for (C1), the old instances for both (C1) and (C2) stay solved.
Thus, in the limit, new problems (instances of problems, more precisely) are created, they are solved in one of the next steps, problems once solved stay solved, so there remain no unsolved problems.
The set of problems increases, but all are solved in the limit, and and are contradictory.
We have (essentially) an alternation of creating and treating problems for (C1) and (C2): We cannot treat e.g. a (C1) problem, bevor it was created by solving a (C2) problem - and conversely.
(Depending on organization, we have some liberty: e.g. solving (C1) creates a (C2) problem, but this needs TWO branches. So we might first treat these two branches ((C1) problems) before going back to the new (C2) problem created by the first new branch.)
2 Background: Graph, Logik
Most of the material in this section is either common knowledge, or was already covered in [Sch23b].
2.1 The Logical Side
On the logics side, we work with propositional formulas, which may, however, be infinite.
We will see that we need infinite formulas (of infinite depth) to have nodes to which we cannot attribute truth values. See Fact 2.4 (page 2.4).
We will work here with disjunctive normal forms, i.e. with formulas of the type where and the are propositional variables or negations thereof - most of the time pure conjunctions of negations of propositional variables.
Fact 2.1
Let where the are propositional variables or negations thereof.
Let - the set of choice functions in the sets where
Then
(Argue semantically with the sets of models and general distributivity and complements of sets.)
Contradictions will be between two formulas only, one a propositional variable, the other the negation of the former.
2.2 The Graph Side
We work with directed, acyclic graphs. They will usually have one root, often called In diagrams, the graphs may grow upward, downward, or sideways, we will say what is meant.
Definition 2.1
Nodes stand for propositional variables.
If a node is not terminal, it has also a propositional formula attached to it, sometimes written with the meaning abbreviated The successors of are the propositional variables occuring in Thus, if and are the only successors of in may be but not
Usually, the are (possibly infinite) conjunctions of propositional variables or (in most cases) their negations, which we write etc. We often indicate the negated variables in the graph with negated arrows, like etc. Thus, usually stands for
Example 2.1
in above definition cannot be replaced by etc, as this example shows (we argue semantically, with the central conflictual construction in Yablo’s paper, Yablo Cell, see Section 1.1 (page 1.1), (2)):
Let so so we have consistency, and it does not work for the construction.
Example 2.2
Consider the basic construction of a contradiction (used by Yablo and here, see Section 1.1 (page 1.1).
stands for so which is impossible.
If we negate then so is possible.
From the graph perspective, we have two paths in from to and
We add now to so thus so and and is not contradictory any more.
Definition 2.2
(This applies only to unique occurrences of a variable in the formula attached to another variable.)
We can attribute a value to a path expressing whether it changes a truth value from the beginning to the end. does not change the value of compared to that of does. We say or positive (negative) path.
Formally:
Let be paths as usual.
If then if then
Let be the concatenation of and Then iff and - otherwise.
If all arrows are negative, then iff the length of is even.
Definition 2.3
We call two paths with common start and end contradictory, and the pair a contradictory cell iff The structures considered here will be built with contradictory cells.
Remark 2.2
Note that the fact that and are contradictory or not is independent of how we start, whether for both or for both
We saw already in Example 2.2 (page 2.2) that it is not sufficient for a “real” contradiction to have two contradictory paths.
We need
(somewhere) an “AND”, so both have to be valid together, an “OR” is not sufficient,
we must not have a branching with an “OR” on the way as in an “escape” path, unless this leads again to a contradiction.
Notation 2.1
When we give nodes a truth value, we will use (and etc. if has the form to denote the case and etc. for the case
Fact 2.3
(Simplified).
Consider three paths, for simplicity with same origin, i.e.
No contradiction loops of length 3.
Suppose they meet at a common point, i.e. Then it is impossible that contradicts contradicts contradicts (at (“ contradicts ” means here that for some but one has value the other value -.)
(Trivial, we have only 2 truth values).
Suppose, first and meet, then (or and meet, but once they meet, they will continue the same way (e.g., if then for all ). Then it is again impossible that contradicts contradicts contradicts and continue to be the same but with different truth values until they meet so it is the same situation as above.)
Above properties generalize to any loops of odd length (with more paths).
This does not hold when the paths may branch again after meeting, as the next Example shows.
Example 2.3
(Example 7.4.2 in [Sch22].)
Let
then contradicts in the lower part, and in the upper part, contradicts and in the upper part, contradicts in the lower part.
Obviously, this may be generalized to paths.
Consider Yablo’s original construction:
Example 2.4
Let the nodes be and the arrows for expressed as a relation by and as a logical formula by
Thus For any we have a contradiction by and for any and for any for a suitable
2.3 Interplay of the Graph and the Logical Side
We can either think on the logical level with formulas, or graphically with conflicting paths, as in the following Fact.
We need infinite depth and width in our constructions:
Fact 2.4
The construction needs infinite depth,
the logic as used in Yablo’s construction is not compact,
the construction needs infinite width, i.e. the logic cannot be classical.
Proof: See Fact 1.4, 858 in [Sch23b].
2.4 A Third Truth Value
See also Section 7.1.2, Fact 7.1.1 in [Sch22] for a less systematic approach.
Remark 2.5
We consider a third truth value, in addition to and
means for a formula that neither nor are true (in the given set of formulas). Of course, this is impossible in a classical framework.
Let be the truth value of
Let be an abbreviation for and
Let
We have
analogously.
Likewise
We thus have:
and
Using the operations for truth values
and above results, we have
(or and thus the order
(and finally the inverse operations
Definition 2.4
We loosely use the expression “escape possibility”, “escape path” etc. to indicate a construction which avoids a desired property through a (usually infinite) sequence of decisions.
Examples:
A simple example is the trivial contradiction, so (See Definition 3.1 (page 3.1).) If we continue with so etc. we have an escape sequence etc., which never meets a contradiction.
Thus, is not a contradiction cell for our purposes, has a model defined by the escape sequence.
“Pipelines”: We may try to avoid infinite branching by one branch used to branch infinite often, but sequentially.
Say, instead of a disjunction of infinitely many arrows we construct one arrow and branch or or etc., but now the sequence …. avoids hitting any
In general, suppose we have a partial graph with origin and end point (among perhaps other end points), and, say results in choices or (due to internal choices by OR).
Remark 2.6
Suppose we have an escape problem, like thus
Can we append a new finite structure to which reduces the possibilities to just one value, say for So, whatever the input, or the outcome is
This, however, is not possible.
There are, whatever the inner structure, just four possibilities: it might be equivalent to The first two cases are trivial. Suppose Then Suppose Then
3 Elementary Cells
3.1 General Remarks
We try to describe here the basic constructions of contradictions.
As said in Fact 2.1 (page 2.1) (and re-written graph-theoretically), these consist of two branches with common origin, which meet again, and have different polarity. We call such constructions cells.
First, we want to exclude some trivialities. We describe only the part beginning at the branching point, not before, not after.
Second, we will define a simple hierarchy of such cells, and will allow a cell only the use of simpler cells as substructure. This prevents “cheating”. (See Remark 3.1 (page 3.1).) Thus, we look at those cells only which allow to construct a Yablo-like structure without the use of more complicate cells.
We first use (almost) only negative arrows, and nodes whose formulas are conjunctions. We will see how to generalize to more complicated paths.
Definition 3.1
We will argue that it suffices to consider these types of cells, see Remark 3.3 (page 3.3) and Section 3.2 (page 3.2). Furthermore, we can also neglect all but the (slightly generalized) Yablo Cells, as we will see in the present section.
It will become clear in a moment that above cells are fundamentally different, but for this we have to consider the second requirement (C2).
Remark 3.1
Suppose we add to the Yablo construction an additional node (or will have the same properties (i.e. truth value as but this addition will not contribute anything to the construction.
To prevent this, we require that it is possible to construct a Yablo structure using only the contradiction cell discussed (and perhaps simpler contradiction cells).
Remark 3.2
We recall again the requirement (C2), has to be contradictory. As (or and thus (or we add the requirement that and (or and are each contradictory.
Thus, Yablo cells and diamonds are different from the simple contradiction.
Yablo cells and diamonds are different from each other as long as we require in the latter case that the truth value of too is and not just False. We will see that making and leads to problems.
(Conversely, making in a Yablo cell not contradictory makes the construction, of course, equivalent to the simple contradiction.)
Remark 3.3
False, True
We present a systematic treatment of variants of the Yablo triangle (and the diamond).
In the Yablo construction, we attach to the Yablo triangle to and constructions, which make We also consider here the cases where we end by or and set e.g. etc., see (2.3.1) below.
We consider:
Do we have a contradiction for
Do we have contradictions for
Do we have escape possibilities?
Do we obtain
- (2.2.5)
We consider the cases in all possible combinations:
Case is equivalent to case
Thus, we look at e.g. case i.e. and etc., and consider whether the requirements in (2.2) above are satisfied.
We first look only at the minimal requirements
has to be contradictory, condition (2.2.1) above.
must not be classical condition (2.2.4) above.
The cases:
Consider so for any so these cases are classical.
Consider For is classical, so is, too. so Thus, is the only non-classical case among
Consider In cases so too. In case so In case so So is the only classical case among the
So the interesting cases left are
We now check for local contradiction in the case
This is simpler, we just have to consider the cases (e.g.) the are omitted, as we look for local contradictions, is anything, and, of course,
The case is known (Yablo construction, all cases except and
Case and we have no contradiction.
Case and we have a contradiction.
So are contradictory.
Next, we have to look at and see if every path from leads to a contradiction, condition (2.2.2) above.
We have to check it must not be
So and will not work, i.e.
Consider i.e. Then in and so this does not work. so this works. so this works.
Consider i.e. Then or and all cases work, except
In summary: in exactly the following cases:
We summarize the three conditions above:
(2.2.4): are ok.
(2.2.2): are not ok, so we learn nothing new beyond (2.2.4).
(2.2.1): are ok, so only is eliminated from those satisfying (2.2.4).
In Summary, only satisfy all three conditions, (2.2.1), (2.2.2), (2.2.4).
We look at these cases.
Here, and in (2.3.6), stands for which is of type likewise as we will append in the full structure at and a construction with value
In all cases
Thus,
Thus,
Thus,
identical to
Thus,
Thus,
We look at the Diamond.
Thus,
The modified Diamond is the same, only so and so we have a Yablo cell.
Summary of the (in this context) important constructions
3.2 Comments on Yablo Cells
Prerequisites:
We consider a triangle , , but leave open if the arrows are positive or negative. We require that is of type and that the triangle is contradictory for the case as well as for both paths und have to lead to a contradiction (conditions (C1) and (C2)).
3.2.1 Branching points
We consider the hierarchy:
intermediate point - branching point - branching point with contradiction
We have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).
If we have a contradiction on only at the situation is again equivalent to the trivial contradiction.
More general, if and contradict each other, and both are for x- valid paths (i.e. no contradiction at then we have an escape possibility.
The points
Suppose we branch at so we have, in addition to the triangle, some
Case 1: If then AND, thus if x- then OR:
Consider , , so, whatever the choice in we have what we need.
Without a contradiction at we have an escape possibility for (see above), e.g. continue the same way: , ,
Case 2: implies OR: (Yablo)
Thus, we also need for a contradiction, this leads to infinite branching at
A new path as in Yablo’s construction leads to infinite branching and non-classical logic.
We may branch on the already existing oder , and continue to build a (finite) contradiction to
This however, is possible only a finite number of times, e.g. then etc., like “Pipeline”, Example 7.37, and Section 7.6.3 in [Sch22] or Example 1.5 and Section 3.4 in [Sch23b]. See also Definition 2.4 (page 2.4) above. This constructs an infinite sequence of choices which offers an escape possibility.
Thus we need infinite branching (and non-classical logic).
We see that we may consider only the new paths they form an infinite subsequence of the construction. By full transitivity of the final construction, any infinite subsequence of the construction has the same properties as the full construction.
3.2.2 Paths
- (1)
We branch on , and have , ,
Case 1: OR:
We make a copy of the structure for
Case 2: AND:
So x- OR:
We make a copy of the structure for
Consider e.g. , ,
As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (” pipeline “), we may assume for simplicity that we always start contradictions at
Diagram 3.1
3.2.3 All Sides Of the Triangle Have to Be Negative
Consider , . Suppose all are of the type
To obtain a contradiction for the triangle needs one or three negative sides.
Assume it has only one negative side.
Let , be a contradiction for (condition (C1)) and , be a contradiction for (for condition (C2)).
Suppose is positive, so we have or and, analogously, or (they have to be of the same type).
We leave aside the condition between and it is only the branching at we are interested in.
Thus, we have or
In the first case so and both and are possibilities, and, as in the simplest contradiction for we have an escape situation (see Definition 2.4 (page 2.4), (1)) (” watered down “by the additional possibilities and
In the second case so and the situation is analogous to the first case.
Suppose is negative, so we have and analogously
Now, we have to contradict for the case but for this we need a third negative arrow so this is impossible, and leads back to the usual Yablo cell.
There is a way out. Instead of considering we continue by This is equivalent to the Yablo cell by
But this is “cheating”.
3.2.4 The Problem with Diamonds
Remark 3.4
We have a conflict between the diamonds starting at and and the diamond starting at
If then y- and As the choices at and are independent, any branch -z, combined with any branch -z, must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
However, if is positive, the diamond -z, has to be contradictory, so not both branches may be negative.
A solution is to “synchronise” the choices at and which can be done e.g. by the formula
and This is a different type of formula, moreoever corresponding arrows from are missing. Even if we do not consider the paths for the diamond starting at we have and so we can make and false, and chose and which results in an escape possibility (we have to consider new diamonds starting at etc.)
This formula has an additional flaw, we want to speak about paths we chose, and not just their end points. This can be done with introducing additional points on the arrows, e.g. the arrow is replaced by now we can speak about the arrow we have given it a name.
We may also put such information in a background theory, which need no be negated.
But all this leads too far from the basically simple formalism of Yablo’s construction, so we will not discuss this any further.
(Yablo’s construction does not need synchronisation, this is done automatically by the universal quantifier.)
The simplified version with recursion only on the left is nothing but the Yablo triangle:
Thus, we conclude that the original version with 4 points and full recursion is beyond our scope, and the simplified version (see above) is nothing but the Yablo triangle.
Diagram 3.2
Nested Diamonds, Details
Diagram 3.3
Diamond Essentials, Recursion on the Left, Simple Contradiction on the Right
4 Combining Cells
4.1 Overview
We follow Yablo’s construction.
(C1): We have to make contradictory.
We first try to make “directly” contradictory. That is, we will not first go to and make contradictory, nor will we make a disjunction at and then make each disjunct contradictory.
(This is not an important simplification, as more complicated constructions will finally use our simple one, too. See e.g. Diagram 4.3 (page 4.3) and its discussion.)
As Yablo did, we set and
The first step is done, is contradictory. Probabilistically, of the possibilities via are contradictory.
(C2) We have to make contradictory.
Now, must be contradictory, too. As we do not know which one is positive, we must make both and contradictory - just as a disjunction is false if every disjunct is false.
Consider, e.g., Again, we will make directly contradictory (as above for so is a conjunction, say (we might re-use here, but this will not be important) and add e.g. and to for the contradiction, so we have
But now, we have destroyed the contradiction at as for is a disjunction, and have to make ALL possibilities for contradictory for so far only the possibility is contradictory.
Probabilistically, of the possibilities at are open.
To create a contradiction for the possibility we introduce new contradictions by adding to so
Etc.
See below and Section 3.3.3. in [Sch23b].
Note that for (C1), i.e. for and for (C2), i.e. for we have to make for any disjunction all possibilities contradictory, and for any conjunction it suffices to make one conjunct contradictory, as and
Of course, following a path in the graph, conjunctions for will become disjunctions for one of the confusing aspects of reasoning.
(C1) and (C2) are antagonistic requirements. Satisfying (C1) creates new problems for (C2), and vice versa. In the limit, all requirements are satisfied, since all problems will solved in the next step, at the price of creating new problems, which will be satisfied in the next step.
Note that we don’t approximate, as long we are finite, we always have only in the infinite case we have
The right level of abstraction:
The author tried first to work towards a representation result using single arrows. However, the following result suggests that one should work with (perhaps composite) paths instead of arrows: a single negative arrow may be replaced by the following diagram:
So
Thus, we may replace negative arrows recursively by arbitrarily deep constructions using negative arrows again, complicating the same construction arbitrarily.
The same is true for all other logical operators.
(See Section 4 in Sch23b.)
4.2 In More Detail
We now consider the inductive construction of the Yablo structure, see Diagram 4.1 (page 4.1) and Diagram 4.2 (page 4.2)
In (2) and (5), we have unrelated points and we may consider the model sets to be orthogonal. - We only indicate model sets briefly, without going into details (U will stand for the universe).
abbreviated
has to lead to a contradiction for
preparation,
was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
finish,
(branch has to be contradicted for add
has to lead to a contradiction for
preparation,
finish,
(branch has to be contradicted for add
(branch has to be contradicted for add
Construction 4.1
Summary:
Thus, there are arrows for all and the construction is transitive for the
Every is head of a Yablo cell with knee
Thus, every arrow from any to any goes to the head of a Yablo cell, and not only the arrows from
This property is “accidental”, and due to the fact that for any arrow there is also an arrow and property (2) holds for by prerequisite.
The construction has infinite depth and branching.
Diagram 4.1
Diagram Inductive Construction 1
Diagram 4.2
Diagram Inductive Construction 2
4.2.1 Composition
The prerequisites for the constructions in the diagrams Diagram 4.1 (page 4.1) etc. are negative Paths, und formulas.
These prerequisites hold also in our generalization to paths (modulo some simplifications for origins at etc.), so these constructions are also valid in the generalization to paths.
Remark 4.1
The condition “Existence of negative paths” is not trivial. Suppose that and are both negative, and is the only path from to then the condition is obviously false. The author does not know how to characterize graphs which satisfy the condition. Some kind of “richness” for the graph will probably have to hold. (This is covered in Yablo’s construction by transitivity.)
Example 4.1
There is a fundamental difference between the contradictions (for and
Consider the case The path is blocked at So we cannot go to by two contradicting possibilities. The path is, however, not blocked at (or elsewhere), so we have two contradicting possibilities to go from to This basically behaves like and we have an escape possibility. So, if we try to append the same construction at we have an escape possibility, always making the origin of the construction negative. This is not true for We can append the same construction at as will always be at the opposite polarity of due to , the alternative path ends at so there is no escape. So a negative start at the original diagram leads to a positive start in the appended diagram.
Diagram 4.3
Diagram Postponed Contradictions
4.3 Beyond Conjunctions
We work with disjunctive normal forms.
Consider
As has to be contradictory, both components and have to be contradictory.
As has to be contradictory, all components, etc., have to be contradictory. For this, it is sufficient and necessary that in each component, one element is contradictory. This is equivalent to the condition that (both and or (both and are contradictory.
Thus
For both
and
have to be contradictory, this is done as usual by
and
For we have to attach a contradiction at either
(-1) and
or
(-2) and
We thus have to examine the combinations
(-1)
(-2)
(-1)
(-2)
The combinations (-1)) and (-2)) are as usual, the other ones are new.
It seems that there might be a problem.
In the cases and seems to be independent from and both solvable by classical logic, a contradiction. But all cases are parts of and and (-1)) and (-2)) need simultaneous treatment of the and the - part.
So the confusion is only transient.
(In the case we have no independent parts, so the “problem” just discussed will not occur.)
5 Conjecture
Consider structures of the type (like ’s), with one small addition: we also allow formulas of the type either as (stand alone), or as one of several elements of a conjunction where is of the type
(This is more general than adding the constant TRUE, as it permits chaining of the new formulas, but we also add TRUE for terminal
In addition, we assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it. We also assume we use only one type of contradiction cell, and work not first with unsuitable cells, postponing the use of efficient cells. This covers also the “interdiction” of escape paths. See here e.g. the beginning of Section 3.1 (page 3.1), and Remark 3.1 (page 3.1), Definition 2.4 (page 2.4), and Example 4.1 (page 4.1), with Diagram 4.3 (page 4.3). Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
5.1 Conjecture: A Suitable Injection Does the Job of Characterisation
Consider Yablo’s original structure YS, and some directed acyclic graph YS’.
Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory iff
There is an injection from YS to YS’, more precisely from the nodes of YS to the nodes of YS’, s.t., if there is a path from to in YS, then there is a path from to
For THIS injection, there is a valuation of the arrows in (the other arrows are not interesting) s.t., if form a contradiction in YS, then likewise for (Contradictions are preserved under
5.2 Idea of Proof (Outline)
Fact 5.2
Yablo’s structure is the simplest structure showing that both and are contradictory.
Proof
Earlier sections of this paper. More precisely:
- (1.1)
- (1.2)
- (1.3)
Remark 5.3
Consider
We may chain such formulas together, like etc., which is impossible for the constant TRUE.
is preserved, if then too. Likewise, if and then
Consider Yablo adding
EG: so , this works well classically, but also for as
We may also continue beyond hinaus: or then too.
If we have mixed chains of and the last element will decide the result (trivial).
Fact 5.4
Adding new formulas inside a Yablo Triangle does not improve anything, it only destroys. It leads either to no contradiction, or to a contradiction of the simplest type, and thus to an escape situation, see Definition 2.4 (page 2.4).
There are no other new contradictory pairs of paths.
Proof
Consider a YC
Insert in ab, or and So, if then if then
Consider now insertions of xy, or xz, yz (analogously) and arbitrary combinations thereof.
xz: If then of then so we have a trivial contradiction and an escape possibility.
xy: xy as so If then if then we have again a trivial contradiction and an escape possibility.
yz: so If then if then we have again a trivial contradiction and an escape possibility.
Combinations:
For combinations of xy and yz, see (2.5) above.
xy and xz: the result is TRUE (both positive) or trivial contradiction with ESC.
xz and yz: This is analogous to xy and xz.
Let be such paths originating in containing some xy as in Fact 5.4 (page 5.4). Then the same considerations as above apply, so we have either no contradiction, or the trivial one with an escape possibility.
Similar results hold when using the constant TRUE: If has -TRUE in one branch, then will have TRUE in this branch, violating (C2). If has TRUE in one branch, then this branch is irrelevant, and we have no contradiction. So TRUE must not occur.
Extend the interpretation beyond
Let If has no successor, then set
If has successors etc., then set
If then has successors. Take the conjunction of the following: if if then
We have to show that still does what it should.
During construction, will be and after completion, And any will be TRUE or By the combination with AND, will not change.