### v4

# Further Comments on Yablo’s Construction ††thanks: File: WAB, [Sch25], arXiv 2504.10370

Abstract
We continue our analysis of Yablo’s coding of the liar paradox by infinite acyclic graphs, [Yab82].
The present notes are based on and continue the author’s previous results on the problem.
In particular, our approach is often more systematic than before.
Contents
1 Introduction
Please note that arXiv pdf files sometimes have problems with some of the diagrams. In that case, please consult the html version.
1.1 Outline
These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions. See also the author’s earlier versions on arXiv. In particular, the reader is referred there for a discussion of more complicated formulas attached to the nodes. We liberally use material from these sources.
All these papers are, of course, based on Yablo’s seminal [Yab82].
In Section 1.2 (page 1.2), we first present the background, its logical and graph theoretical side, and their interplay. We also discuss a useful third truth value, which expresses that neither nor is consistent.
In most cases, we will work with simple conjunctions (with and without negations).
In Section 2 (page 2), we discuss elementary contradictions (cells). Essentially, these are variations of Yablo’s triangles. The simplest contradictions are too simple, they have an “escape possibility”, i.e., combining them will leave one possibility without contradictions. Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework.
In Section 3 (page 3), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction.
We then generalize slightly, see Remark 3.1 (page 3.1), but show that this general construction contains Yablo’s construction, so it is minimal. (See also the discussion of “Saw Blades” in Section 7.6 of [Sch22] and Section 3 in [Sch23b].)
In both cases, we show that - within the chosen variant - the construction is necassary.
- (4)
There is an important property for the graphs, which we need, probably a sort of “richness”, see Remark 1.5 (page 1.5). This property holds trivially in Yablo’s construction by transitivity. In our approach, it will hold by the construction of a suitable injection, see Section 4 (page 4), Condition 4.1 (page 4.1) (2).
1.2 Basic Definitions and Facts
1.2.1 Terminology
Yablo’s construction
Yablo’s construction is an infinite directed graph with nodes and (negative) arrows for
Nodes stand for propositional variables, and the meaning of arrows is that variable occurs (negatively) in the formula attached to in most cases it expresses (We will identify for simplicity nodes with their attached variables.)
It is easy to see that no node may be true, nor false: Suppose is true, than all are false, so must be false, then contradiction, as all must be false. Suppose is false, then some must be true, again a contradiction (to the above).
Yablo Cell, Yablo Triangle
The basic contradiction structure has the form with the meaning (so and
We sometimes call the head, the knee, and the foot of the Yablo Cell (or Triangle).
Conditions (C1) and (C2)
We sometimes write when we want to emphasize that we assume is positive, similarly for the negative case.
Consider in the Yablo construction (but this applies to all
We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction. (More precisely, see Section 3.1 (page 3.1), (3).)
1.2.2 The Logical Side
On the logics side, we work with propositional formulas, which may, however, be infinite.
We will see that we need infinite formulas (of infinite depth) to have nodes to which we cannot attribute truth values. See Fact 1.6 (page 1.6).
Notation 1.1
When we give nodes a truth value, we will use (and etc. if has the form to denote the case and etc. for the case
We will work here with disjunctive normal forms, i.e. with formulas of the type where and the are propositional variables or negations thereof - most of the time pure conjunctions of negations of propositional variables.
Fact 1.1
Let where the are propositional variables or negations thereof.
Let - the set of choice functions in the sets where
Then
(Argue semantically with the sets of models and general distributivity and complements of sets.)
Contradictions will be between two formulas only, one a propositional variable, the other the negation of the former.
Fact 1.2
Let and we want to show that is contradictory, then it suffices to show that a subset of the is inconsistent - this corresponds to the fact that (X any truth value.)
Let and we want to show that is contradictory, then we have to show that ALL are contradictory. Again, this corresponds to
The following example illustrates the propagation of required contradictions.
Example 1.1
Consider
As has to be contradictory, both components and have to be contradictory.
As has to be contradictory, all components, etc., have to be contradictory. For this, it is sufficient and necessary that in each component, one element is contradictory. This is equivalent to the condition that (both and or (both and are contradictory.
Thus
For both
and
have to be contradictory, this is done as usual by
and
For we have to attach a contradiction at either
(-1) and
or
(-2) and
We thus have to examine the combinations
(-1)
(-2)
(-1)
(-2)
The combinations (-1)) and (-2)) are as usual, the other ones are new.
It seems that there might be a problem.
In the cases and seems to be independent from and both solvable by classical logic, a contradiction. But all cases are parts of and and (-1)) and (-2)) need simultaneous treatment of the and the - part.
So the confusion is only transient.
(In the case we have no independent parts, so the “problem” just discussed will not occur.)
1.2.3 The Graph Side
We work with directed, acyclic graphs. They will usually have one root, often called In diagrams, the graphs may grow upward, downward, or sideways, we will say what is meant.
Definition 1.1
Nodes stand for propositional variables.
If a node is not terminal, it has also a propositional formula attached to it, sometimes written with the meaning abbreviated The successors of are the propositional variables occuring in Thus, if and are the only successors of in may be but not
Usually, the are (possibly infinite) conjunctions of propositional variables or (in most cases) their negations, which we write etc. We often indicate the negated variables in the graph with negated arrows, like etc. Thus, usually stands for more precisely.
Example 1.2
in above definition cannot be replaced by etc, as this example shows (we argue semantically, with the central conflictual construction in Yablo’s paper, Yablo Cell, see Section 1.2.1 (page 1.2.1), (2)):
Let so so we have consistency, and it does not work for the construction.
Example 1.3
Consider the basic construction of a contradiction (used by Yablo and here, defined “officially” in Definition 2.1 (page 2.1).
stands for so which is impossible.
If we negate then so is possible.
From the graph perspective, we have two paths in from to and
We add now to so thus so and and is not contradictory any more.
Definition 1.2
(This applies only to unique occurrences of a variable in the formula attached to another variable.)
We can attribute a value to a path expressing whether it changes a truth value from the beginning to the end. does not change the value of compared to that of does. We say or positive (negative) path.
Formally:
Let be paths as usual.
If then if then
Let be the concatenation of and Then iff and - otherwise.
If all arrows are negative, then iff the length of is even.
Definition 1.3
We call two paths with common start and end contradictory, and the pair a contradictory cell iff The structures considered here will be built with contradictory cells.
Remark 1.3
Note that the fact that and are contradictory or not is independent of how we start, whether for both or for both
We saw already in Example 1.3 (page 1.3) that it is not sufficient for a “real” contradiction to have two contradictory paths.
We need
(somewhere) an “AND”, so both have to be valid together, an “OR” is not sufficient,
we must not have a branching with an “OR” on the way as in an “escape” path, unless this leads again to a contradiction.
Fact 1.4
(Simplified).
Consider three paths, for simplicity with same origin, i.e.
No contradiction loops of length 3.
Suppose they meet at a common point, i.e. Then it is impossible that contradicts contradicts contradicts (at (“ contradicts ” means here that for some but one has value the other value -.)
(Trivial, we have only 2 truth values).
Suppose, first and meet, then (or and meet, but once they meet, they will continue the same way (e.g., if then for all ). Then it is again impossible that contradicts contradicts contradicts and continue to be the same but with different truth values until they meet so it is the same situation as above.)
Above properties generalize to any loops of odd length (with more paths).
This does not hold when the paths may branch again after meeting, as the next Example shows.
Example 1.4
(Example 7.4.2 in [Sch22].)
Let
then contradicts in the lower part, and in the upper part, contradicts and in the upper part, contradicts in the lower part.
Obviously, this may be generalized to paths.
Consider Yablo’s original construction:
Example 1.5
Let the nodes be and the arrows for expressed as a relation by and as a logical formula by
Thus For any we have a contradiction by and for any and for any for a suitable
Example 1.6
This Example shows that infinitely many finitely branching points cannot always replace infinite branching - there is an infinite “procrastination branch” or “escape branch”. This modification of the Yablo structure has one acceptable valuation for
Let as usual, and introduce new
Let with
If then by so, generally,
if then and
If then so if etc., so, generally,
if then or
Suppose now then for all and for all By there is and a contradiction, or for all again a contradiction.
But is possible, by setting and for all
Thus, replacing infinite branching by an infinite number of finite branching does not work for the Yablo construction, as we can always chose the “procrastinating” branch.
Diagram 1.1
Escape Path
Remark 1.5
For a representation result, we need enough paths in the graph considered to form contradictions. This is not trivial. Suppose that and are both negative, and is the only path from to then the condition is obviously false. The author does not know how to characterize graphs which satisfy the condition directly. Some kind of “richness” for the graph will probably have to hold. (This is covered in Yablo’s construction by transitivity.)
1.2.4 Interplay of the Graph and the Logical Side
We can either think on the logical level with formulas, or graphically with conflicting paths, as in the following Fact.
We need infinite depth and width in our constructions:
Fact 1.6
The construction needs infinite depth,
the logic as used in Yablo’s construction is not compact,
the construction needs infinite width, i.e. the logic cannot be classical.
Proof
Let be a minimal element, then we can chose an arbitrary truth value and propagate this value upwards. If there are no infinite descending chains, we can do this for the whole construction.
The logic as used in Yablo’s construction is not compact:
Trivial. (Take This is obviously inconsistent, but no finite subset is.).
It is impossible to construct a Yablo-like structure with classical logic:
Take an acyclic graph, and interpret it as in Yablo’s construction. Wlog., we may assume the graph is connected. Suppose it shows that cannot be given a truth value. Then the set of formulas showing this does not have a model, so it is inconsistent. If the formulas were classical, it would have a finite, inconsistent subset, Define the depth of a formula as the shortest path from to this formula. There is a (finite) such that all formulas in have depth Give all formulas of depth (arbitrary) truth values, and work upwards using truth functions. As the graph is acyclic, this is possible. Finally, has a truth value.
Thus, we need the infinite
The following example illustrates (partially) unfinished contradictions, and resulting escape possibilities.
Example 1.7
There is a fundamental difference between the contradictions (for and
Consider the case The path is blocked at So we cannot go to by two contradicting possibilities. The path is, however, not blocked at (or elsewhere), so we have two contradicting possibilities to go from to This basically behaves like and we have an escape possibility. So, if we try to append the same construction at we have an escape possibility, always making the origin of the construction negative. This is not true for We can append the same construction at as will always be at the opposite polarity of due to , the alternative path ends at so there is no escape. So a negative start at the original diagram leads to a positive start in the appended diagram.
Diagram 1.2
Diagram Postponed Contradictions
1.2.5 A Third Truth Value
See also Section 7.1.2, Fact 7.1.1 in [Sch22] for a less systematic approach.
Remark 1.7
We consider a third truth value, in addition to and
means for a formula that neither nor are true (in the given set of formulas). Of course, this is impossible in a classical framework.
Let be the truth value of
Let be an abbreviation for and
Let
We have
analogously.
Likewise
We thus have:
and
Using the operations for truth values
and above results, we have
(or and thus the order
(and finally the inverse operations
1.3 Strategy
These comments will be clearer in hindsight, after having read the details. Still the best place to put them seems to be in general remarks, and go into details later on.
The comments on “ is false ” show that there is no other way to express this in an acyclic graph of our type.
The comments on the construction in the limit solve a seeming contradiction by chosing the right level of abstraction, i.e. by arguing on the level of instances.
1.3.1 Expressing “x is false” In an Acyclic Graph
In an acyclic graph, we cannot express something like we have to use different means to express that is impossible.
As we can always fill in truth values (including from the bottom, we cannot force by looking upwards of (or sideways, i.e. there is no monotone path connecting with some So we have to look downward.
We may use or if we have the constant We will see later, that this not interesting, see Definition 2.1 (page 2.1) and Remark 2.1 (page 2.1). This is a contradiction made of (graphically speaking) two paths of length one each. Interesting contradictions are made of slightly more complicated paths, like the Yablo Cell meaning
Basically, the only way to express that is impossible, is by the logical formula
A slightly different argument: How can we express that is wrong in a finite acyclic graph? Consider the set of minimal elements below , and the formula corresponding to the graph in DNF of these elements. This might be
(the graph is trivial, just
or a tautology or contradiction.
a disjunction of conjunctions of elements in , or their negations. By acyclicity,
This argument is not totally satisfactory. For this reason, we reformulate:
We examine graphs, where there existe a node such that as well as -x will lead to a contradiction.
1.3.2 Satisfying the Antagonistic Requirements (C1) and (C2) In the Limit
We want to make and contradictory. We know that this is impossible in finite graphs, so we have to show how we solve the problem in the (infinite) limit. We proceed inductively as Yablo did, but we also see that we basically have no other choice (modulo minor modifications).
So we want to express by an infinite graph that both and are inconsistent. As we saw in Section 1.3.1 (page 1.3.1), this must be expressed by suitable contradictions. Therefore, we treat such contradictions as basic building blocks, as Yablo did.
The problems are then
- (1)
How do we combine them to obtain a suitable graph expressing that both and are impossible?
The problem is that, if is contradictory, then is a tautology in classical logic, and thus in the theory expressed by a finite acyclic graph. So we alternate between and a tautology, and it is not clear what such constructions are in the limit, when we leave classical logic.
This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 3 (page 3).
A summary: we have to look at instances.
- (2.1)
- (2.2)
But is now a branching point. The branch continues to be satisfied for (C1), but the (new) branch opens a new problem for (C1), as is an OR for
- (2.4)
So we have new instances of problems for (C1), (C2), we solve the new instances, say for (C1), the old instances for both (C1) and (C2) stay solved.
Thus, in the limit, new problems (instances of problems, more precisely) are created, they are solved in one of the next steps, problems once solved stay solved, so there remain no unsolved problems.
The set of problems increases, but all are solved in the limit, and and are contradictory.
We have (essentially) an alternation of creating and treating problems for (C1) and (C2): We cannot treat e.g. a (C1) problem, bevor it was created by solving a (C2) problem - and conversely.
(Depending on organization, we have some liberty: e.g. solving (C1) creates a (C2) problem, but this needs TWO branches. So we might first treat these two branches ((C1) problems) before going back to the new (C2) problem created by the first new branch.)
1.4 The “Right” Level of Abstraction
1.4.1 Introduction
The basic elements in Yablo’s construction are negative arrows, from which Yablo Cells are built. We show here that we can build arbitrarily complex structures equivalent to negative arrows, and as a matter of fact to any propositional logical operator. This suggests that the right level of abstraction to consider more general Yablo-like structures is not the level of single arrows, but rather of paths.
Remark 1.8
In the diagrams in Diagram 1.3 (page 1.3) the inner path is barred by , and the outer path contains 3 or 2 negations.
Definition 1.4
We define negations and their type.
A negation (diagram) is a diagram all of whose arrows are of the kind.
We define negation diagrams, or, simply, negations, and their type.
An arrow is a negation of type 0.
all negations inside are of type and at least one negation inside is of type
Remark 1.9
Thus, it seems very difficult to describe Yablo type diagrams on the level of single arrows. It seems to be the wrong level of abstraction.
We now give some examples.
1.4.2 Some Examples
Example 1.8
Negation (1)
so and
Negation (2)
so and continue as above.
Identity
so and
TRUE
Thus, and
so and
Diagram 1.3
Propositional Formulas 1
Diagram 1.4
Propositional Formulas 2
1.5 Prerequisites
We work with pure conjunctions of positive or negative propositional variables, like except for the following:
we may use a constant TRUE or
we may use formulas of the type
so the following are also permitted: etc.
This is used only to “neutralize” unwanted parts of the graph we characterize, see Fact 4.4 (page 4.4).
- (2)
- (3)
- (4)
2 Elementary Cells
2.1 Introduction
Elementary cells have to be simple constructions providing contradictions (in some cases, e.g. when the origin, usually called or so, cannot have a certain truth value, as this will lead, within the cell, to a contradiction).
Thus
They must be contradictory, and thus exclude at least one instance.
In addition, they should be simple, e.g. they will not contain a sequence or so without branching, as we could replace this with
(We will see later how to relinquish this requirement.)
More importantly, we must be able to construct a Yablo-like structure, built up by this type of elementary cell only (and perhaps some simpler auxiliary structures). They must serve as building blocks for a suitable stucture. This prevents “cheating”, letting other constructions do the crucial work.
- (4)
2.2 Basic Discussion
See Section 7.3 in [Sch22] for more material.
We try to describe here the basic constructions of contradictions.
As said in Fact 1.1 (page 1.1) (and re-written graph-theoretically), these consist of two branches with common origin, which meet again, and have different polarity. We call such constructions cells.
First, we want to exclude some trivialities. We describe only the part beginning at the branching point, not before, not after.
Second, we will define a simple hierarchy of such cells, and will allow a cell only the use of simpler cells as substructure. This prevents “cheating”. (See Section 2.1 (page 2.1).) Thus, we look at those cells only which allow to construct a Yablo-like structure without the use of more complicate cells.
We first use (almost) only negative arrows, and nodes whose formulas are conjunctions. We will see how to generalize to more complicated paths.
Definition 2.1
We will argue that it suffices to consider these types of cells, see Remark 2.3 (page 2.3) and Section 2.4 (page 2.4). Furthermore, we can also neglect all but the (slightly generalized) Yablo Cells, as we will see in the present section.
It will become clear in a moment that above cells are fundamentally different, but for this we have to consider the second requirement (C2).
Remark 2.1
An escape problem with for the trivial contradiction, so (See Definition 2.1 (page 2.1).) If we continue with so etc. we have an escape sequence etc., which never meets a contradiction.
Thus, is not a contradiction cell for our purposes, has a model defined by the escape sequence.
Suppose we have an escape problem, like thus
Can we append a new finite structure to which reduces the possibilities to just one value, say for So, whatever the input, or the outcome is
This, however, is not possible.
There are, whatever the inner structure, just four possibilities: it might be equivalent to The first two cases are trivial. Suppose Then Suppose Then
In general, suppose we have a partial graph with origin and end point (among perhaps other end points), and, say results in choices or (due to internal choices by OR).
Example 2.1
We consider now some simple, contradictory cells. They should not only be contradictory for the case but also be a potential start for the case without using more complex cells. (Otherwise, we postpone the solution, and may forget the overly simple start.)
We will consider in Section 2.5 (page 2.5) a more complicated contradiction cell (“Diamond”), and discuss here only simple cells.
Note that the following considerations apply also to cells formed by paths consisting of more than one arrow, and not only to cells formed by single arrows. See Section 2.4 (page 2.4). In particular, also in more complicated cells all sides have to negative.
The cell with 2 arrows.
It corresponds to the formula graphically, it has a positive and a negative arrow from to so exactly one of and is negative.
If is positive, we have a contradiction.
Of course, appending at a Yablo Cell (see below, case (2.3)) may be the beginning of a contradictory structure, but this is “cheating”, we use a more complex cell.
Cells with 3 arrows.
In the graph with not annotated arrows, they have the form -z, -z, where the lines may stand for or
Again, we want a contradiction for positive, so we need an at For the same reason, the number of negative arrows may not be even. So assume the number of negative arrows is one or three.
Three negative arrows:
This is the original type of contradiction in Yablo’s construction
This will be discussed in detail in the rest of the paper. But we see already that both paths, and change sign, so and will be positive, appending the same type of cell at and solves the problem (locally), and offers no escape.
We call sometimes the head of the Yablo cell, its foot, and its knee.
If we combine Yablo Cells, the knee for one cell may become the head for another Cell, etc.
One negative arrow:
We will show that all cases have an escape possibility, along the positive arrow (or arrows) originating in So they are not suitable for a construction using only this type of cell.
Consider the case
Upper part:
Lower part:
Consider the case
Consider the case
Here,
Diagram 2.1
Simple Contradictions
Diagram 2.2
Diagram ESC-1
Diagram 2.3
Diagram ESC-2
Diagram 2.4
Diagram ESC-3
Remark 2.2
The distinction between and i.e. between knee and foot, is conceptually important. In the case we have at a complete contradiction, at we have not yet constructed a contradiction. Thus, if we have at again an (as at this becomes an and we have to construct a contradiction for all (or not only for Otherwise, we have an escape possibility for Obviously, the contradiction need not be immediate at the important property is that ALL paths through all lead to a contradiction, and the simplest way is to have the contradiction immediately at - as in Yablo’s construction, and our saw blades.
See also the discussion of “Saw Blades” in [Sch22], Section 3.
As we work for a contradiction in the case x-, too, the simplest way to achieve this is to have a negative arrow and at again an This gives a chance to construct a contradiction at Of course, we have to construct a contradiction at too, as in the case we have an at
Of course, we may have branches originating at which all lead to contradictions in the case so (resp. at is possible, too.
But, for the simple construction, we need between and and at And this leads to the construction of contradictions for all (or as just mentioned above.
2.3 Values for y and z in Yablo Cells
We consider here what happens when we give and values In addition, we consider the variants and
Remark 2.3
False, True
We present a systematic treatment of variants of the “environment” of the Yablo triangle (and the diamond).
In the Yablo construction, we attach to the Yablo triangle to and constructions, which make We also consider here the cases where we end by or and set e.g. etc., see (3.1) below.
Note that will always hold.
We consider the following requirements:
Do we have a contradiction for i.e. will not be true in the Yablo Cell?
Will both branches lead to possible contradictions for
Do we obtain
- (2.4)
We consider the cases in all possible combinations:
Case is equivalent to case
Thus, we look at e.g. case i.e. and etc., and consider whether the requirements in (2) above are satisfied.
We first look only at the minimal requirements
has to be contradictory, requirement (2.1) above.
must not be classical (i.e. not or requirement (2.3) above.
The cases:
(2) (requirement (2.3))
Consider so for any so these cases are classical.
Consider For is classical, so is, too.
so
Thus, is the only non-classical case among
Consider In cases so too.
In case so
In case so
So is the only classical case among the
So the interesting cases left are
(1) (requirement (2.1))
We now check for local contradiction in the case i.e. must not be
may only be iff But if then only in case
Next, we have to look at and see if every path from leads to a contradiction, requirement (2.2) above. So neither nor must be
So and will not work.
Consider i.e.
Then in and so this does not work.
so this works.
so this works.
Consider i.e.
Then or and all cases work, except
In summary: in exactly the following cases: so they are excluded.
We summarize the results for the three requirements above:
(2.3): are ok.
(2.2): are not ok, so we learn nothing new beyond requirement (2.3).
(2.1): Only is eliminated, so we learn nothing new beyond requirement (2.3).
In summary, only satisfy all three requirements, (2.1), (2.2), (2.3).
We look at these cases.
Here, and in (3.6), stands for which is of type likewise as we will append in the full structure at and a construction with value
In all cases
Thus,
Thus,
Thus,
identical to
Thus,
Thus,
We look at the Diamond.
Thus,
The modified Diamond is the same, only so and so we have a Yablo cell.
Summary of the (in this context) important constructions
2.4 Yablo Cells Formed By Paths
We consider a triangle , , where etc. may be composite paths. We know from Section 2.2 (page 2.2) that the paths have to be negative, but they may be composed of positive and negative arrows.
We require that is of type and that the triangle is contradictory for the case as well as for both paths und have to lead to a contradiction (conditions (C1) and (C2)). Note that and have to be the first nodes on above paths to be contradictory for - otherwise we will not reach (or the path is barred before.
2.4.1 Branching points
We have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).
Branching at
Suppose we branch at so we have, in addition to the triangle, some
We know that is negative, so, if then This is the situation of the Yablo construction, see Section 3 (page 3). So just a short comment here.
Thus, we also need for a contradiction, this leads to infinite branching at
A new path as in Yablo’s construction leads to infinite branching and non-classical logic.
We may branch on the already existing oder , and continue to build a (finite) contradiction to
This however, is possible only a finite number of times, e.g. then etc., see Example 1.6 (page 1.6). This constructs an infinite sequence of choices which offers an escape possibility.
Thus we need infinite branching (and non-classical logic).
We see that we may consider only the new paths they form an infinite subsequence of the construction. By full transitivity of the final construction, any infinite subsequence of the construction has the same properties as the full construction.
2.4.2 Paths
Additional branching on , e.g. , ,
Case 1: OR
This generates a copy of the structure for
We construct for as for , i.e.
Case 2: AND
So OR, so we make a copy of the structure for consider e.g. , ,
Note: We simplify, suppose e.g. that we have then we do not need another contradiction at the one at suffices (condition (C1)). We need, however, that all branches lead to contradictions (condition (C2)).
We branch on , and have , .
Case 1: OR:
We make a copy of the structure for
Case 2: AND:
So x- OR:
We make a copy of the structure for
Consider e.g. , , .
As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (” pipeline “), we may assume for simplicity that we always start contradictions at
Diagram 2.5
2.5 The Problem with Diamonds
Remark 2.4
We have a conflict between the diamonds starting at and and the diamond starting at
If then y- and As the choices at and are independent, any branch -z, combined with any branch -z, must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
However, if is positive, the diamond -z, has to be contradictory, so not both branches may be negative.
A solution is to “synchronise” the choices at and which can be done e.g. by the formula
and This is a different type of formula, moreoever corresponding arrows from are missing. Even if we do not consider the paths for the diamond starting at we have and so we can make and false, and chose and which results in an escape possibility (we have to consider new diamonds starting at etc.)
This formula has an additional flaw, we want to speak about paths we chose, and not just their end points. This can be done with introducing additional points on the arrows, e.g. the arrow is replaced by now we can speak about the arrow we have given it a name.
We may also put such information in a background theory, which need no be negated.
But all this leads too far from the basically simple formalism of Yablo’s construction, so we will not discuss this any further.
(Yablo’s construction does not need synchronisation, this is done automatically by the universal quantifier.)
The simplified version with recursion on one side only, here on the left only, is nothing but the Yablo triangle:
Thus, we conclude that the original version with 4 points and full recursion is beyond our scope, and the simplified version (see above) is nothing but the Yablo triangle.
Diagram 2.6
Nested Diamonds, Details
Diagram 2.7
Diamond Essentials, Recursion on the Left, Simple Contradiction on the Right
3 Combining Cells
3.1 Overview
We follow Yablo’s construction.
(C1): We have to make contradictory.
We first try to make “directly” contradictory. That is, we will not first go to and make contradictory, nor will we make a disjunction at and then make each disjunct contradictory.
(This is not an important simplification, as more complicated constructions will finally use our simple one, too. See e.g. Diagram 1.2 (page 1.2) and its discussion.)
As Yablo did, we set and
The first step is done, is contradictory. Probabilistically, of the possibilities via are contradictory.
(C2) We have to make contradictory.
Now, must be contradictory, too. As we do not know which one is positive, we must make both and contradictory - just as a disjunction is false if every disjunct is false.
Consider, e.g., Again, we will make directly contradictory (as above for so is a conjunction, say (we might re-use here, but this will not be important) and add e.g. and to for the contradiction, so we have
But now, we have destroyed the contradiction at as for is a disjunction, and have to make ALL possibilities for contradictory for so far only the possibility is contradictory.
Probabilistically, of the possibilities at are open.
To create a contradiction for the possibility we introduce new contradictions by adding to so
Etc.
See below and Section 3.3.3. in [Sch23b].
Note that for (C1), i.e. for and for (C2), i.e. for we have to make for any disjunction all possibilities contradictory, and for any conjunction it suffices to make one conjunct contradictory, as and
Of course, following a path in the graph, conjunctions for will become disjunctions for one of the confusing aspects of reasoning.
(C1) and (C2) are antagonistic requirements. Satisfying (C1) creates new problems for (C2), and vice versa. In the limit, all requirements are satisfied, since all problems will solved in the next step, at the price of creating new problems, which will be satisfied in the next step.
Note that we don’t approximate, as long we are finite, we always have only in the infinite case we have
3.2 In More Detail
We now consider the inductive construction of the Yablo structure, see Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2)
In (2) and (5), we have unrelated points and we may consider the model sets to be orthogonal. - We only indicate model sets briefly, without going into details (U will stand for the universe).
abbreviated
has to lead to a contradiction for
preparation,
was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
finish,
(branch has to be contradicted for add
has to lead to a contradiction for
preparation,
finish,
(branch has to be contradicted for add
(branch has to be contradicted for add
Construction 3.1
Summary:
Thus, there are arrows for all and the construction is transitive for the
Every is head of a Yablo cell with knee
Thus, every arrow from any to any goes to the head of a Yablo cell, and not only the arrows from
This property is “accidental”, and due to the fact that for any arrow there is also an arrow and property (2) holds for by prerequisite.
The construction has infinite depth and branching.
Remark 3.1
Our construction is somewhat arbitrary. We construct a contradiction for by and use the arrow as part of the contradiction for This, of course, is a special case.
Consider Diagram 3.3 (page 3.3). Here, the contradiction for is as above, but we begin a totally new contradiction for likewise for etc.
But, we see, the old construction of Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2) is part of the more general construction of Diagram 3.3 (page 3.3), so the old construction is really minimal.
We have to be a bit careful here. If we set etc. then we have etc. Instead, we may set so we have as desired, or we append at recursively the same construction - as in “Saw Blades”.
Diagram 3.1
Diagram Inductive Construction 1
Diagram 3.2
Diagram Inductive Construction 2
Diagram 3.3
Diagram Inductive Construction Core
4 Result
Consider structures of the type (like ’s), with one small addition: we also allow formulas of the type either as (stand alone), or as one of several elements of a conjunction where is of the type It is important that these additions also cooperate with our truth value
(This is more general than adding the constant TRUE, as it permits chaining of the new formulas, but we also add TRUE for terminal nodes.)
In addition, we assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it. We also assume we use only one type of contradiction cell, and work not first with unsuitable cells, postponing the use of efficient cells. This covers also the “interdiction” of escape paths. See here e.g. the beginning of Section 2.2 (page 2.2), and Remark 2.1 (page 2.1), and Example 1.7 (page 1.7), with Diagram 1.2 (page 1.2). Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
4.1 A Suitable Injection Does the Job of Characterisation
Consider Yablo’s original structure YS, and some directed acyclic graph YS’.
Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory iff the following conditions hold.
Condition 4.1
There is an injection from YS to YS’, more precisely from the nodes of YS to the nodes of YS’, s.t., if there is a path from to in YS, then there is a path from to
For THIS injection, there is a valuation of the arrows in (the other arrows are not interesting) s.t., if form a contradiction in YS, then likewise for (Contradictions are preserved under
4.2 Proof
Fact 4.2
Yablo’s structure is the simplest structure showing that both and are contradictory.
Proof
Earlier sections of this paper. More precisely:
- (1)
- (2)
- (3)
Remark 4.3
Consider
We may chain such formulas together, like etc., which is impossible for the constant TRUE.
is preserved, if then too. Likewise, if and then
Consider Yablo adding
EG: so , this works well classically, but also for as
We may also continue beyond : or then too.
If we have mixed chains of and the last element will decide the result (trivial).
Fact 4.4
Consider (without the other elements etc. in
Now consider the full YS’, and attribute values to the nodes as follows, respecting the values already given in (1):
If is minimal, then
Otherwise:
If dann
arrow
Note that these arrows have no value, i.e. we have no distinction between and
If then
: arrow
: arrow
: arrow
Note that arrows inside have a value, these values have to be preserved. The other arrows have no value, and we treat them as positive.
Moreover, we did not distinguish between minimal and not minimal this does not matter. If we want to distinguish we have to differentiate the last conjunct further.
- (3)
Remark 4.5
Asssume for simplicity that (arrows may be negative, too), likewise for etc.
We have paths etc., assume for simplicity that the only new intermediate node is respectively).
Let is not in
We set arbitrary, Thus, so and so
Of course, instead of going through the additional path might also go through etc. This changes nothing. We have neutralised the path not in
Diagram 4.1
Diagram
Diagram 4.2
Diagram