### v5

# Further Comments on Yablo’s Construction ††thanks: File: WAB, [Sch25], arXiv 2504.10370

Abstract
We show that the existence of a node with attached formula in an acyclic graph such that neither nor its negation are consistent is equivalent to the existence of subgraph equivalent to Yablo’s construction.
Contents
1 Introduction
1.1 Outline
These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions. See also [Sch25]. (Please note that arXiv pdf files sometimes have problems with some of the diagrams. In that case, please consult the html version.) We liberally use material from these sources. All these papers are, of course, based on Yablo’s seminal [Yab82].
Prerequisites:
All graphs considered will by acyclic.
- (1.2)
For ease of exposition, we introduce a third truth value, as abbreviation for “ and are inconsistent”.
We will use only one type of contradictory cell in a graph. This prevents “cheating”.
Contradictions in a graph will begin immediately, and not after some arbitrary path. This simplifies constructions, and is not an important restriction.
- (1.6)
In Section 1.2 (page 1.2), we first present the background, its logical and graph theoretical side, and their interplay. We also discuss a useful third truth value, which expresses that neither nor is consistent. In Section 1.3 (page 1.3), we discuss our strategy, and discuss the antagonistic conditions (C1) and (C2) and how to satisfy them in an inductive construction. (We need infinite depth and width in the construction, see e.g. [Sch25], Fact 2.4, or [Sch23b], Fact 1.4).
In Section 2 (page 2), we discuss elementary contradictions (cells). Essentially, these are variations of Yablo’s triangles. The trivial contradiction, see Definition 2.1 (page 2.1), is too simple, it has an “escape possibility”, i.e., combining them will leave one possibility without contradictions. See Definition 1.6 (page 1.6). Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework. See Section 2.5 (page 2.5). We also show that all sides of the triangles have to be negative, see Remark 2.2 (page 2.2). Section 2.3 (page 2.3) discusses further details of these triangles. We generalize to triangles formed by paths in Section 2.4 (page 2.4).
In Section 3 (page 3), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction. We then generalize slightly, see Remark 3.1 (page 3.1), but show that this general construction contains Yablo’s construction, so it is minimal. This is perhaps the central result of the paper. This fact is essentially re-used in Section 4 (page 4). (See also the discussion of “Saw Blades” in Section 7.6 of [Sch22] and Section 3 in [Sch23b].) In both cases, we show that the construction is necessary.
In Section 4 (page 4), we show that
a graph without valuation can be given a valuation such that there is an element such that neither nor are consistent, if there is a suitable injection from Yablo’s structure into and a partial valuation for the nodes of the paths of the injection, such that contradictions are preserved. We complete the partial injection by a trivial extension.
conversely, if there is a graph with (complete) valuation with some node such that neither nor are consistent, then we can construct a subgraph which corresponds again to an injection of the Yablo structure. We use here again the minimality of Yablo’s construction, this time imside the given graph.
1.2 Basic Definitions and Facts
Notation 1.1
We consider directed, acyclic graphs. They will usually have one root, often called Nodes and arrows may be labelled.
If the nodes are labelled, we identify nodes with their labels, unless said otherwise. The labels of nodes stand for propositional variables or constants The propositional variables will be unique in a given graph. Moreoever, to is associated a formula the direct successors of a node determine the propositional variables or constants occurring in The theory of a graph is the union of its a node in the graph. We sometimes write for the theory of a graph
We sometimes write if occurs negatively in
In many cases is a conjunction of positive or negative occurrences of the direct successors of If not, we will say so expressly. For instance, the graph will express that we often write instead of
Definition 1.1
The pure graph, i.e. just nodes and arrows, tells only which variables occur in the associated formulas We call the annotations or valuations, the pure graph graph without valuations or annotations, and the full graph annotated graph or graph with valuations, etc. If is the pure graph, will sometimes be an annotated version of and conversely.
Definition 1.2
Let be a node in a fixed graph as above.
We say that is contradictory (relative to iff the assumption leads to a contradiction in Likewise for is contradictory, iff the assumption is contradictory in this theory.
Truth values:
We set iff is contradictory, and iff not, likewise for
We introduce a third truth value which is an abbreviation for and - relative to of course. See Remark 1.2 (page 1.2).
When the context is clear, we sometimes simplify to etc.
Contradictions via paths - we will change between the graph and the logical formulation, whatever seems intuitively better:
Such contradictions have in a graph the following form: there is a node and a node and two paths from to which branch at and join again at and they contradict each other, i.e. one branch results in the other in
These are the basic contradictions.
- (3.2)
Consider again Diagram 3.1 (page 3.1), part (2.1). The whole diagram is not contradictory, as we opened a new possibility at which becomes impossible only in part (3) of this diagram. So, as a detail, the paths are contradictory, but the possibility is closed only in (3), so the contradiction of the whole diagram is only complete in (3). If you prefer to think probabilistically, in part (2.1) of the cases are contradictory etc. See also Section 1.3.1 (page 1.3.1) for opening and closing new possibilities.
Definition 1.3
(Yablo, see [Yab82].)
We call the following graph the Yablo structure, YS:
Written as graph: We use negated arrows, and positive arrows, in graphs as this improves readability considerably (see Notation 1.1 (page 1.1).)
Written as set of propositional formulas:
Contradictions in YS:
Suppose is a node in the YS. Then and which is contradictory.
But starting with will also result in a contradiction:
suppose is true, then is contradictory, as we just saw, so is contradictory, too.
Note that we abused notation here, where may be an arrow in a graph, or an implication in a logical formula. Context will tell.
The contradictions have the form this is the basic contradiction “cell”, we call such cells Yablo cells, ’s, or Yablo triangles.
We sometimes call the head, the knee, the foot of such ’s. This is relative to the cell discussed. When we combine such cells, the knee or foot of one cell may become the head of the next cell, etc.
Remark 1.1
Let where the are propositional variables or negations thereof, thus, is in disjunctive normal form, DNF.
Let - the set of choice functions in the sets where
Then
(Argue semantically with the sets of models and general distributivity and complements of sets.)
Example:
Consider the Yablo cell with each formula associated with of the type We then have 16 possible combinations like etc., 7 are in 9 in is in the 9 combinations of and the 9 combinations of By is in the 7 positive combinations of so we have again a contradiction, as usual.
(Alternatively, we may simplify, and work instead of with many combinations, e.g. just with the totally positive combination for and with the totally negative combination
The generalization to arbitrary DNF’s is trivial.
Thus, working with DNF’s changes nothing compared to the Yablo construction, working with All we need is a suitable negation.
Remark 1.2
The truth value has the following properties:
(and, of course
the following order is compatibel with and i.e. for any truth values and
This completely describes the behaviour of
This is the only 3-valued logic with the property and compatibility with and It is thus identical to the Strong Kleene logic Ks, as was pointed out by a referee.
Proof:
For more details, see [Sch25]
trivial
Consider two copies of ’s, YS and YS’, beginning with and respectively, and the following two structures (a) with (b) with
Consider a new origin, and arrows from to as follows: and and
Details are left to the reader.
Easy.
Definition 1.4
(This applies only to unique occurrences of a variable in the formula attached to another variable.)
We can attribute a value to a path expressing whether it changes a truth value from the beginning to the end. does not change the value of compared to that of does. We say or positive (negative) path.
Formally:
Let be paths as usual.
If then if then
Let be the concatenation of and Then iff and - otherwise.
1.3 Strategy
These comments will be clearer in hindsight, after having read the details. Still the best place to put them seems to be in general remarks, and go into details later on.
Definition 1.5
Conditions (C1) and (C2)
We sometimes write when we want to emphasize that we assume is positive, similarly for the negative case.
Consider in the Yablo construction (but this applies to all
We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction.
1.3.1 Satisfying the Antagonistic Requirements (C1) and (C2) In the Limit
We want to make and contradictory. We know that this is impossible in finite graphs, so we have to show how we solve the problem in the (infinite) limit. We proceed inductively as Yablo did, but we also see that we basically have no other choice (modulo minor modifications).
So we want to express by an infinite graph that both and are inconsistent. See Definition 1.2 (page 1.2). Therefore, we treat such contradictions as basic building blocks, as Yablo did.
The problems are then
- (1)
How do we combine them to obtain a suitable graph expressing that both and are impossible?
The problem is that, if is contradictory, then is a tautology in classical logic, and thus in the theory expressed by a finite acyclic graph. So we alternate between and a tautology, and it is not clear what such constructions are in the limit, when we leave classical logic.
This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 3 (page 3), we have to look at instances.
- (2.1)
- (2.2)
But is now a branching point. The branch continues to be satisfied for (C1), but the (new) branch opens a new problem for (C1), as is an OR for
- (2.4)
So we have new instances of problems for (C1), (C2), we solve the new instances, say for (C1), the old instances for both (C1) and (C2) stay solved.
Thus, in the limit, new problems (instances of problems, more precisely) are created, they are solved in one of the next steps, problems once solved stay solved, so there remain no unsolved problems.
The set of problems increases, but all are solved in the limit, and and are contradictory.
We have (essentially) an alternation of creating and treating problems for (C1) and (C2): We cannot treat e.g. a (C1) problem, bevor it was created by solving a (C2) problem - and conversely.
(Depending on organization, we have some liberty: e.g. solving (C1) creates a (C2) problem, but this needs TWO branches. So we might first treat these two branches ((C1) problems) before going back to the new (C2) problem created by the first new branch.)
Example 1.1
This Example shows that infinitely many finitely branching points cannot always replace infinite branching - there is an infinite “escape branch” or path. See below, Definition 1.6 (page 1.6). This modification of the Yablo structure has one acceptable valuation for
Let as usual, and introduce new
Let with
If then by so, generally,
if then and
If then so if etc., so, generally,
if then or
Suppose now then for all and for all By there is and a contradiction, or for all again a contradiction.
But is possible, by setting and for all
Thus, replacing infinite branching by an infinite number of finite branching does not work for the Yablo construction, as we can always chose the “escape” branch, see below.
Definition 1.6
When we try to make a structure contradictory for some value of the origin, but find a way (a valuation) to avoid or escape a contradiction, we call such a valuation an escape path, branch, or possibilty. This is then a model consistent with the theory of the graph.
This seems an intuitive description how the structure fails to be contradictory for this valuation.
2 Elementary Cells
2.1 Introduction
We try to generalize Yablo cells to some other types of cells.
(C1) They will have the form of two paths which branch say at and meet again at (but not before), and are contradictory.
(C2) Both branches will have to lead to new contradiction cells if
We require that it is possible to construct a structure with such cells starting, say at such that and are contradictory.
In addition, we require that this is possible using only the contradiction cells of this type. This prevents “cheating”, where the real work is done by some other type of cell.
We will impose syntactical limitations on the formulas attached to the nodes in the cells.
2.2 Basic Discussion
Definition 2.1
We will argue that it suffices to consider these types of cells, see Remark 2.3 (page 2.3) and Section 2.4 (page 2.4). Furthermore, we can also neglect all but the (slightly generalized) Yablo Cells, as we will see in the present section.
It will become clear in a moment that above cells are fundamentally different, but for this we have to consider the second requirement (C2).
Remark 2.1
- (1)
Suppose we have an escape problem, like thus
Can we append a new finite structure to which reduces the possibilities to just one value, say for So, whatever the input, or the outcome is
This, however, is not possible.
There are, whatever the inner structure, just four possibilities: it might be equivalent to
The first two cases are trivial, as we have the possibilities or so is always a possibility.
Suppose Then
Suppose Then
In general, suppose we have a partial graph with origin and end point (among perhaps other end points), and, say results in choices or (due to internal choices by OR).
If we continue with a copy of now grafting on etc., we may not achieve a desired result (whereas a different may achieve it).
Remark 2.2
We consider here cells with three arrows.
Note that the following considerations apply also to cells formed by paths consisting of more than one arrow, and not only to cells formed by single arrows. See Section 2.4 (page 2.4). In particular, also in more complicated cells all sides have to negative.
In the graph with not annotated arrows, they have the form , , where the lines may stand for or
Again, we want a contradiction for positive, so we need an at For the same reason, the number of negative arrows may not be even. So assume the number of negative arrows is one or three.
Three negative arrows:
This is the original type of contradiction in Yablo’s construction
This will be discussed in detail in the rest of the paper. But we see already that both paths, and change sign, so and will be positive when is negative, appending the same type of cell at and solves the problem (locally), and offers no escape.
One negative arrow:
We will show that all cases have an escape possibility, along the positive arrow (or arrows) originating in So they are not suitable for a construction using only this type of cell.
Consider the case
Consider the theory indicates an escape path.
Consider the theory indicates an escape path.
Consider the case
Consider the theory indicates an escape path.
Consider the case
Consider the theory indicates an escape path.
See also the discussion of “Saw Blades” in [Sch22], Section 3.
2.3 Values for y and z in Yablo Cells
We consider here what happens when we give and truth values In addition, we consider the variants and
Remark 2.3
We present a systematic treatment of variants of the “environment” of the Yablo triangle (and the diamond).
In the Yablo construction, we attach to the Yablo triangle to and constructions, which make We also consider here the cases where we end by or and set e.g. etc., see (3.1) below.
Note that will always hold.
We consider the following requirements:
Do we have a contradiction for i.e. will not be true in the Yablo Cell?
Will both branches lead to possible contradictions for
Do we obtain
Do we obtain infinite width and depth by suitable combinations of cells of the same type, see Section 2.1 (page 2.1), and [Sch25], Fact 2.4 or [Sch23b], Fact 1.4. See also Section 3 (page 3). This requires continuation at both and with constructions of value case See below (4.3) for an (unimportant) exception.
We consider the cases in all possible combinations:
Case is equivalent to case
Thus, we look at e.g. case i.e. and etc., and consider whether the requirements in (2) above are satisfied.
We look at above requirements in the following order: (2.3), (2.1), (2.2). This is a tedious exercise and left to the reader, these are the results:
(2.3): are ok.
(2.2): are not ok, so we learn nothing new beyond requirement (2.3).
(2.1): Only is eliminated, so we learn nothing new beyond requirement (2.3).
In summary, only satisfy all three requirements, (2.1), (2.2), (2.3).
We look at the results in the interesting cases.
Here, and in (3.6), stands for which is of type likewise as we will append in the full structure at and a construction with value
In all cases
Thus,
Thus,
so
(Yablo) Thus,
identical to Thus,
Thus,
We look at the Diamond.
Thus,
The modified Diamond is the same, only so and so we have a Yablo cell.
Summary of the (in this context) important constructions
- (4.1)
- (4.2)
The simplified Saw Blade tooth is of the type with type above, see Section 7.6 in [Sch22], or Section 3 in [Sch23b]. (In this construction, will be (except for the start) again of type as we obtain this by transitivity, forced by condition (C1), see Construction 3.1 (page 3.1), and Diagram 3.1 (page 3.1), steps (2.3), (2.5), (2.6)). Both the original Saw Blade and the simplified one contain the Yablo construction.
See, e.g., Diagram 3.1, upper part, in [Sch23b]: has the role of in Yablo’s construction. is a Yablo cell. must be the head of a new Yablo cell, starting e.g. with has to be contradicted, so we need so must be the head of a new Yablo cell, so we have some etc.
- (4.4)
2.4 Yablo Cells Formed By Paths
We consider a triangle , , where etc. may be composite paths. We know from Section 2.2 (page 2.2) that the paths have to be negative, but they may be composed of positive and negative arrows.
We require that is of type and that the triangle is contradictory for the case as well as for both paths und have to lead to a contradiction (conditions (C1) and (C2)). Note that and have to be the first nodes on above paths to be contradictory for - otherwise we will not reach (or the path is barred before.
2.4.1 Branching points
Remark 2.4
is the origin of a new cell, so we have to branch at (Anyway, we have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).)
Branching at
Suppose we branch at so we have, in addition to the triangle, some
We know that is negative, so, if then This is the situation of the Yablo construction, see Section 3 (page 3). So just a short comment here.
Thus, we also need for a contradiction, this leads to infinite branching at
A new path as in Yablo’s construction leads to infinite branching and non-classical logic.
We may branch on the already existing oder , and continue to build a (finite) contradiction to
This however, is possible only a finite number of times, e.g. then etc., see Example 1.1 (page 1.1). This constructs an infinite sequence of choices which offers an escape possibility.
Thus we need infinite branching (and non-classical logic).
We see that we may consider only the new paths they form an infinite subsequence of the construction. By full transitivity of the final construction, any infinite subsequence of the construction has the same properties as the full construction.
Branching at on the path :
Consider the case Whatever the branching situation on is (just one branching, or multiple branchings), for the branch considered in , the path is supposed to create a contradiction. Suppose we have a branching at where i.e. it is an “OR”, then we may decide at otherwise, and not have a contradiction with the branch . Thus, we first chose a path via and then chose a contradicting path (or directly ). In Yablo’s construction, this is trivial, as is a direct link.
Definition 2.2
If , then there is forming a Yablo cell, i.e. all three components , , are negative, such that contains no branching point with for
We call this weak transitivity.
2.4.2 Paths
Additional branching on , e.g. , ,
Case 1: OR
This generates a copy of the structure for
We construct for as for , i.e.
Note: need not be contradictory for , as is contradictory, thus, will be contradictory ( will be AND for ), so (C2) is satisfied. Thus, has the role of and we may invoke our rule that contradictions should happen immediately, and not be postponed.
Case 2: AND
So OR, so we make a copy of the structure for consider e.g. , ,
Note: We simplify, suppose e.g. that we have then we do not need another contradiction at the one at suffices (condition (C1)). We need, however, that all branches lead to contradictions (condition (C2)).
We might identify and , and impose this as an additional requirement for the structures we examine, along with “no postponement of contradictions”.
We branch on , and have , .
Case 1: OR:
We make a copy of the structure for
Case 2: AND:
So x- OR:
We make a copy of the structure for
Consider e.g. , , .
As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (“pipeline”), we may assume for simplicity that we always start contradictions at
2.5 The Problem with Diamonds
Remark 2.6
We have a conflict between the diamonds starting at and and the diamond starting at
If then and . As the choices at and are independent, any branch , , combined with any branch , , must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
However, if is positive, the diamond , has to be contradictory, so not both branches may be negative, likewise for
A solution is to “synchronise” the choices at and Informally, this is possible by: “ and at choose the arrows or and at choose the arrows or ”.
However we express this, the negation is “ or choose at or choose at the arrow ”, which has nothing to do with the usual negations in Yablo structures. A way out might be to put the choices at and into a background theory which is not affected by negation.
Even if we consider arbitrary DNF formulas, we cannot speak about choices at and or arrows like Thus, within our framework, the diamond is NOT a suitable contradiction cell.
(Yablo’s construction does not need synchronisation, this is done automatically by the universal quantifier.)
The simplified version with recursion on one side only, here on the left only, is nothing but the Yablo triangle:
Diagram 2.1
Nested Diamonds, Details
3 Combining Cells
We follow Yablo’s construction, and refer the reader to Section 1.3.1 (page 1.3.1) for an abstract discussion of this construction.
Construction 3.1
The principle
If has to be the start of a Yablo cell, but is a terminal node, add an arrow
If has to be the start of a Yablo cell, exists, but not add (transitivity).
If will be the start of a new Yablo cell, if
Details
thus
will be the start of new Yablo cells.
By (C2), has to lead to a contradiction for
preparation: we use as part of the new Yablo cell, and add the new arrow as second part of the new cell, where is a new point.
finish: We add the new arrow as third part of the new cell.
Note: was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
By (C1),
(branch has to be contradicted for add thus
will be the start of a new Yablo cell.
By (C2), has to lead to a contradiction for
preparation: we use as part of the new Yablo cell, and add the new arrow as second part of the new cell, where is a new point.
finish: We add the new arrow as third part of the new cell.
By (C1),
(branch has to be contradicted for add Thus
Complete Yablo cell starting at
By (C1),
(branch has to be contradicted for add Thus
will be the start of a new Yablo cell.
Note:
Summary (in the limit):
Main properties:
Thus, there are arrows for all and the construction is transitive for the
Every is head of a Yablo cell with knee
Thus, every arrow from any to any goes to the head of a Yablo cell, and not only the arrows from
This property is “accidental”, and due to the fact that for any arrow there is also an arrow and (C2) holds for by prerequisite.
The construction has infinite depth and branching.
Remark 3.1
Our construction is somewhat arbitrary. We construct a contradiction for by and use the arrow as part of the contradiction for This, of course, is a special case.
We may begin a totally new contradiction for likewise for etc.
But the old construction of Diagram 3.1 (page 3.1) is part of this somewhat more general construction, so the old construction is really minimal.
We have to be a bit careful here. If we set etc. then we have etc. Instead, we may set so we have as desired, or we append at recursively the same construction - as in “Saw Blades”.
Remark 3.2
Above construction can be read differently: Suppose we have an (annotated) graph such that a node has truth value Then we can do above construction. E.g., has to be contradictory, (C1), so have two contradictory paths from to, say and some on the path Both paths have to lead to contradictions for (C2). Exactly one of them has to be on one of the paths (and not at the end), say this is (if both are on the paths, we would have the diamond). This way, we can reconstruct the Yablo construction within Of course, there may be additional branchings, as discussed in Section 2.4 (page 2.4). Next, we show how to neutralise them.
Definition 3.1
Consider an annotated graph a subset of its nodes, and the subgraph generated by in Consider now where in is identical to in restricted to the nodes in
Remark 3.3
above is identical up to logical equivalence for the nodes in to the following modification of replace for all and arbitrary, if and is terminal by and if is not terminal, by We call tis the isolation of in
Example: Let be as follows: and then
Consequently, in Remark 3.2 (page 3.2), the isolation is equivalent to the original Yablo structure.
Diagram 3.1
Diagram Inductive Construction
4 Result
4.1 Prerequisites (Limitations)
We assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it. We also assume we use only one type of contradiction cell, and work not first with unsuitable cells, postponing the use of efficient cells. This covers also the “interdiction” of escape paths. See here e.g. the beginning of Section 2.2 (page 2.2), and Remark 2.1 (page 2.1). Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
Given these limitations, we show that Yablo’s structure is minimal for acyclic structures showing that and lead to contradictions, and that any unlabelled, directed graph permits a labelling showing the same iff Yablo’s original structure has a suitable injection into the graph in question.
4.2 A Suitable Injection From the Yablo Structure
Consider Yablo’s original structure YS, and some directed acyclic graph YS’.
Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory if the following conditions hold.
Condition 4.1
There is an injection from YS to YS’, more precisely from the nodes of YS to the nodes of YS’, s.t., if there is a path from to in YS, then there is a path from to in YS’, in particular, if is a direct successor of on such a path, there must be an arrow in YS’.
For this injection, we need the following condition
Let be the subgraph consisting of all nodes and the intermediate nodes on these paths.
We have to be able to construct a partial valuation in YS’, defined on the nodes in YS” (of course respecting the arrows in s.t. for any two nodes if there is an arrow in YS, then iff This need not hold for intermediate nodes. Thus, contradictions are preserved for the paths in YS’.
Definition 4.1
We extend this partial valuation in YS’ to a total valuation in YS’ as follows:
If is a node in set there is an arrow in Thus, we respect the arrows in YS’, and assures that the result is and not if
If is a node in YS”: We have already a formula describing the behaviour for all successors of in YS”. We now have to respect the other arrows in YS’, and set
there is an arrow in YS’, and not a successor of in above paths
Remark 4.1
- (1)
The second condition expresses “richness”, that we have enough independent arrows.
4.3 For the converse
Fact 4.3
Yablo’s structure is the simplest structure showing that both and are contradictory, and contained in all such structures.
Proof
Earlier sections of this paper. More precisely:
- (1)
- (2)
- (3)
Fact 4.4
Suppose the graph has a valuation with an element
Proof
By Fact 4.3 (page 4.3), or directly by the construction Construction 3.1 (page 3.1) INSIDE we have an analogue of the Yablo structure in perhaps with paths instead of simple arrows, and perhaps with many unnecessary parts.
We can eliminate the unnecessary parts as follows, (for ease of reading with a simple example, analogously to Definition 4.1 (page 4.1)):
Suppose we have where the part is unnecessary. Then replace by so we have
Suppose we have where is unnecessary, then replace by so we have
As we have DNF formulas, this takes care of it.
5 Acknowledgements
The author would like to thank the anonymous referee for his many valuable suggestions, and the editorial board and Ms. Jane Spurr for their help.